OpenStudy (anonymous):

Solve -7 < x + 2 < 4. A. -9 < x < 2 B. -5 < x < 6 C. x < 2 or x > -9 D. no solution

OpenStudy (kropot72):

If you subtract 2 from each of the 3 terms what do you get?

OpenStudy (anonymous):

D?

OpenStudy (kropot72):

No. What is the result of each of these; -7 - 2 = ? x + 2 -2 = ? 4 - 2 = ?

OpenStudy (anonymous):

oh, I accidentally did -2 - 7 lol

OpenStudy (anonymous):

A?

OpenStudy (kropot72):

A is the right choice. Well done :) BTW: -2 - 7 = -9 and -7 - 2 = -9. The order doesn't matter in this case.